OpenStudy (anonymous):

The legs of a right triangles measure (m − 3) in. and (n + 3) in. Which is the length of square of its hypotenuse?

OpenStudy (anonymous):

m2 + n2 − 6(m − n) + 18 m2 + n2 + 6(m − n) + 18 m2 + n2 + 6(m − n) + 9 m2 + n2 − 6(m − n) ?

OpenStudy (anonymous):

1st. choice

OpenStudy (anonymous):

Alright, thank you!

OpenStudy (anonymous):

Using the Pythagorean Theorem: \[ \mbox{the square of the hypotenuse}= (m-3)^2+ (n+3)^2\] \[= m^2 -3m-3m+9+n^2+3n+3n+9 \] \[=m^2+n^2-6m+6n+18 \] \[=m^2+n^2-6(m-n)+18\]